A market research company finds that a the price of p=$20, they would sell x=42 tiles each month. If they lower the price to p=$10, then more people would purchase the tiles, and they can expect to sell x=52 tiles in a month's time. find the equation of the line for the demand equation. write the answer in the form p=mx +b. Hint: write an equation using points in the form (x,p).
Answer by ewatrrr(24785)


Hi
(42,20)
(52,10) m = -10/10 = -1
p = -x + b |using ordered pair (42,20) to solve for x
20 = -42 + b
62 = b
p = -x + 62